The components of a force are 5N and 12N along east and north directions ,the forces has magnitude are
1
Expert's answer
2020-09-14T10:19:38-0400

The magnitude is


"F=\sqrt{F_x^2+F_y^2}=\sqrt{5^2+12^2}=13\ N"